a light ray is incident on a transparent sphere of index root 2 at an angle of incidence =45 o . what is the deviation of a tiny fraction of ray which enters sphere undergoes two internal reflections and then refracts out into air

Grade:12

1 Answers

Vikas TU
5 years ago
At first surface apply the Snell's law, 
n1*sint(thetha1) = n2*sin(thetha2)
root2*sinthetha2 = 1/root(2) 
thetha2 = 30 degree.
The refrcting angle that is r = 30 degree.
 
Also from TIR concept, r = r' = 30° 
Apply the Snell's law again at second surface 
we get,
e = 45° 
Deviation at the first surface
= i - r = 45° - 30° = 15° 
Deviation at 2nd surface 
= e - r = 45° - 30° = 15° 
Total Deviation = 15° + 15° = 30°
(total deviateion always gets add up)